OpenStudy (anonymous):

1. (–22 + 17) + 13 = –22 + (17 + 13)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

thanks for helping

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

how do you know

OpenStudy (anonymous):

Because they are the same equation on each side.

OpenStudy (anonymous):

ok what about –(27 + (–3)) = –27 + 3

OpenStudy (anonymous):

Communitive.

TheSmartOne (thesmartone):

@Aleah.wood0613 DO no give direct answers!

TheSmartOne (thesmartone):

It is against the CoC http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

35 + (–15 + 4) = (35 + (–15)) + 4

TheSmartOne (thesmartone):

And @shadybubba Make a new question and post it there.

OpenStudy (anonymous):

dont do it.

TheSmartOne (thesmartone):

You are not advised to post 15 questions in 1 post.

OpenStudy (anonymous):

im not

OpenStudy (anonymous):

get out b###

OpenStudy (anonymous):

35 + (–15 + 4) = (35 + (–15)) + 4

TheSmartOne (thesmartone):

Is this a test @shadybubba @~ ??

OpenStudy (anonymous):

what about this 35 + (–15 + 4) = (35 + (–15)) + 4

OpenStudy (anonymous):

associative

TheSmartOne (thesmartone):

@ganeshie8

OpenStudy (anonymous):

and last one –12 + 6 = 6 + (–12)

TheSmartOne (thesmartone):

@Zarkon Cheating..

OpenStudy (anonymous):

Diss.

TheSmartOne (thesmartone):

@Abhisar @ganeshie8